OpenStudy (anonymous):

(x+4) over 3 = 2 is the answer x=2 ? (:

OpenStudy (hba):

Yes.

OpenStudy (anonymous):

thanks!(:

OpenStudy (hba):

You can always check your answer anytime without posting it lol :P Lemme show you x+4/=2 2+4/3=2 6/3=2 2=2 LHS=RHS Hence your answer is correct :P You're welcome ;)

OpenStudy (hba):

@sylinan Got it ?

OpenStudy (anonymous):

kinda (:

OpenStudy (hba):

No worries lol c: